Benzene-1,2,3-tricarboxylic acid (C6H3 (CO2H) 3) is a three-protic, water-soluble carboxylic acid with the following pKa values: pKa1 = 2.86, pka2 = 4.30 and pka3 = 6.28. What will be the pH of a 0.075 M solution of this treprotic carboxylic acid?
Answer #1

1 (A+B+1) - 10 6.28 he kag = V (A+B+1) =) -6.78 EN (A+B+V) (B-1 (B-1) (B-1)zo Salung . and get [x] = 9.50822x103M [B] =4.9601pH of the solution has been calculated.